• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kiadtisak

    Log entry time 15:22:10 on June16,2010

    Entry number 326925

    keyword=Database updated

    Databases for left and right arms staring on 04/14/2010 have been updated after taking the beam position scan into account. Starting from 06/04/2010, a new db_L.vdc.dat has been changed due to T0 error. The right arm already looks good with the same database.
    I've attached two sieve slit plots for run 27727 and 6841. Compared to Bob's halog 324421, the new database improves the sieve plots. However, the right arm looks much better than the left arm.


    FIGURE 1

    FIGURE 2